OpenStudy (anonymous):

someone help me pleasee medal and fan ! Write the equation for the vertical line that contains the point (1,1) a. y=0x+1 b. y=x c. y=1 d. x=1

OpenStudy (anonymous):

the answer is x=1 cause it stays at 1 x and y always changes

OpenStudy (anonymous):

Write the equation of the line passing thought the points (-2,7) and (5,-7) in slope-intercept form a. y=-2x-3 b.y=-2x+3 c. y=2x-3 d. y=2x+3

OpenStudy (anonymous):

@marinesrock

OpenStudy (anonymous):

The first thing I'd do is find the slope, m, between these two points, can you do that? 

OpenStudy (anonymous):

i think its b

OpenStudy (rajat97):

the x co-ordinate is equal to y so the answer is b) x=y or else, you can put the values x=1 and y=1 in any of the options